OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at (7, 0) and a radius of 10?

OpenStudy (anonymous):

(x -h)^2 + (y-k)^2 = r2 center is (h,k)

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

uw

OpenStudy (anonymous):

I think that's center-radius form... some texts call that standard form. General form is usually an equation with one side of the equal sign being zero. So to get general form from center radius form, foil out your center-radius form and make one side of the equal sign zero.

OpenStudy (anonymous):

General Formula: X^ 2 + Y^ 2=r^2 where r is the radius, with ccenter at 0,0 since your center is not at 0,0 - you must use the one that I provided earlier.

OpenStudy (anonymous):

The general form for the equation of a circle is Ax^2 + Ay^2 + Cx + Dy + E = 0 where A, C, D, and E are real.